OpenStudy (anonymous):

Let S be the set of numbers of the form n(n + 1)(n + 2)(n + 3)(n + 4), where n is any positive integer. The first few terms of S are: 1*2*3*4*5 = 120 2*3*4*5*6= 720 3*4*5*6*7= 2520 and so no . What is the GCD of the elements of S?

OpenStudy (anonymous):

Greatest common factor for s, i think it will be the numbers used for the 5 times table

OpenStudy (anonymous):

emmm.. if I put this in form n(n + 1)(n + 2)(n + 3)(n + 4) = S and 5 times the table that is to say what should be the value of n?, it's easy but the neck of answering or interpretation ....am not able to understand

OpenStudy (mathmath333):

\(\large \color{black}{\begin{align}k=n(n+1)(n+2)(n+3)(n+4)\end{align}}\) \(\large \color{black}{\begin{align} \begin{array}{|c|c|c|} \hline \text{for n=1} & k=120 &k=2^3\times 3\times 5\\ \hline \text{for n=2} & k=720&k=2^4\times 3^2\times 5\\ \hline \text{for n=3} & k=2520&k= 2^3\times 3^2\times 5\times 7\\ \hline \text{for} n=\cdots & k=\cdots&k=\cdots\\ \hline \end{array} \end{align}}\) from the table \(\large \color{black}{\begin{align}\gcd\left(n(n+1)(n+2)(n+3)(n+4)\right)=2^3\times 3\times 5=120 \end{align}}\)

ganeshie8 (ganeshie8):

product of any \(n\) consecutive integers is divisible by \(n!\)

OpenStudy (mathmath333):

120 is wrong ?

ganeshie8 (ganeshie8):

120 is right !

OpenStudy (welshfella):

I'm a bit confused about GCD and GCF What is the difference?

OpenStudy (mathmath333):

\(\large \color{black}{\begin{align}\gcd=\text{gcf}= \text{hcf}\end{align}}\)

OpenStudy (mathmath333):

http://en.wikipedia.org/wiki/Greatest_common_divisor

OpenStudy (mathmath333):

http://mathworld.wolfram.com/GreatestCommonDivisor.html

OpenStudy (welshfella):

thanks - basically they are the same thing

ganeshie8 (ganeshie8):

factor = divisor yes they are exact same thing

OpenStudy (anonymous):

thanks you guys are real fast in solving @ganishie8 and @ mathmath333

OpenStudy (welshfella):

I was thinking of LCD where D means the denominator

ganeshie8 (ganeshie8):

\[b=ac\] \(a\) is a `divisor` of \(b\) \(a\) is a `factor` of \(b\) \(b\) is a `multiple` of \(a\)

ganeshie8 (ganeshie8):

also notice \(0\) cannot be a divisor of any natural number because \(x= 0y\) is possible only when \(x=y=0\)

OpenStudy (mathmath333):

cool!